# Universal bounds on optimization of free energy harvesting

### 20 Dec 2023 v3

@@ Acknowledgements.
Acknowledgements.— We thank members of the Complex Systems Lab, Bernat Corominas-Murtra, Luis Seoane, David Wolpert, and Damian Sowinski for useful discussions. AK also thanks Sosuke Ito for support and encouragement. JP was partly supported by María de Maeztú fellowship MDM-2014-0370-17-2. This project has received funding from the European Union’s Horizon 2020 research and innovation programme under the Marie Skłodowska-Curie Grant Agreement No. 101068029. It also received support from Grant 62417 from the John Templeton Foundation. The opinions expressed in this publication are those of the authors and do not necessarily reflect the views of the John Templeton Foundation.
@@
Universal bounds on optimization of free energy harvesting
@@
Derivations of main results
1.1   Derivation of (4) from LDB and steady-state assumption
@@
$\dot{\mathcal{G}}^{\mathrm{tot}}=\dot{\mathcal{F}}^{b}(\bm{\pi})+\dot{\mathcal{G}}^{b}(\bm{\pi})-\dot{\Sigma}({\boldsymbol{J}^{c}})/\beta.$ (S1)
@@
$={\beta}^{-1}\sum_{i,j}p_{i}R^{c}_{ji}[\ln p_{j}+\beta(f_{j}+{g}^{c}_{ji})].$ (S2)
The first line follows from definitions made in the main text. Since $R^{c}$ is a rate matrix, $\sum_{i,j}p_{i}R^{c}_{ji}a_{i}=\sum_{i}p_{i}a_{i}\sum_{j}R^{c}_{ji}=0$ for any $\bm{a}$. This allows us to further rewrite (S2) as
@@
$\ln(R^{c}_{ji}/R^{c}_{ij})=\beta(f_{i}-f_{j}-{g}^{c}_{ji})\qquad\text{for}\quad R^{c}_{ji}>0.$ (S4)
Using the definition of one-way fluxes due to control given a generic distribution $\bm{p}$, ${J^{c}_{ji}}=p_{i}R^{c}_{ji}$, and the Schnackenberg expression of EPR $\dot{\Sigma}$, we rewrite (S3) as
˙$\dot{\mathcal{F}}^{c}(\bm{p})+\dot{\mathcal{G}}^{c}(\bm{p})=-{\beta}^{-1}\sum_{i,j}p_{i}R^{c}_{ji}\ln\frac{p_{j}R^{c}_{ji}}{p_{i}R^{c}_{ij}}=-{\beta}^{-1}\sum_{i\neq j}{J^{c}_{ji}}\ln\frac{{J^{c}_{ji}}}{{J^{c}_{ij}}}=-{\beta}^{-1}\dot{\Sigma}({\boldsymbol{J}^{c}})\,.$ (S5)
In the main text, this equality is applied at the steady-state distribution $\bm{\pi}$. However, Eq. (S5) is valid for all distributions $\bm{p}$ as long as (S4) holds.
To complete our derivation of (S1), recall that $R\bm{\pi}=(R^{b}+R^{c})\bm{\pi}=0$ by definition of the steady-state $\bm{\pi}$. At the same time, from the definition of ˙$\dot{\mathcal{F}}^{b}(\bm{p})$ in (1), we have
˙$\dot{\mathcal{F}}^{b}(\bm{\pi})+\dot{\mathcal{F}}^{c}(\bm{\pi})=\sum_{j}(f_{j}+{\beta}^{-1}\ln\pi_{j})\sum_{i}(R^{b}_{ji}+R^{c}_{ji})\pi_{i}=0\qquad\implies\qquad\dot{\mathcal{F}}^{b}(\bm{\pi})=-\dot{\mathcal{F}}^{c}(\bm{\pi}).$ (S6)
Combining the definition of $\dot{\mathcal{G}}^{\mathrm{tot}}$ with (S5) and (S6) gives Eq. (S1),
$\dot{\mathcal{G}}^{\mathrm{tot}}=\dot{\mathcal{G}}^{b}(\bm{\pi})+\dot{\mathcal{G}}^{c}(\bm{\pi})=\dot{\mathcal{G}}^{b}(\bm{\pi})-\left[\dot{\mathcal{F}}^{c}(\bm{\pi})+{\beta}^{-1}\dot{\Sigma}({\boldsymbol{J}^{c}})\right]=\dot{\mathcal{F}}^{b}(\bm{\pi})+\dot{\mathcal{G}}^{b}(\bm{\pi})-{\beta}^{-1}\dot{\Sigma}({\boldsymbol{J}^{c}})\,.$ (S7)
@@
$\dot{\Sigma}({\boldsymbol{J}^{c}})=\frac{1}{2}\sum_{i\neq j}(p_{j}R^{c}_{ij}-p_{i}R^{c}_{ji})\ln\frac{p_{j}R^{c}_{ij}}{p_{i}R^{c}_{ji}}\geq 0,$ (S8)
@@
Second, recall our assumption that the control processes do not interact directly with the external source of free energy, but only with the internal reservoir and heat bath. This assumption is formalized in the particular form of the LDB condition (S4). It means that for the control processes, the combined “system+internal reservoir” may be treated as a single system coupled only to a heat bath at inverse temperature $\beta$. Hence, the rate of entropy production due to control is simply $\beta$ times the decrease of the combined free energy of the system and internal reservoir, as in (S5). Moreover, owing to Second Law, their combined free energy cannot increase under control dynamics [24, 33],
˙$\dot{\mathcal{F}}^{c}(\bm{p})+\dot{\mathcal{G}}^{c}(\bm{p})\leq 0\qquad\text{for all}\quad\bm{p}.$ (S9)
1.2   Concavity of constrained optimization (5)
Consider the objective ˙$\dot{\mathcal{F}}^{b}(\bm{p})+\dot{\mathcal{G}}^{b}(\bm{p})-\dot{\Sigma}({\boldsymbol{J}})/\beta$ in (5) in the main text. In Sec. 1.3 below, we show that the term ˙$\dot{\mathcal{F}}^{b}(\bm{p})+\dot{\mathcal{G}}^{b}(\bm{p})$ is concave in $\bm{p}$. Here we show that $\dot{\Sigma}({\boldsymbol{J}})$ is convex in ${\boldsymbol{J}}$. This shows that the overall objective is concave in the pair $(\bm{p},{\boldsymbol{J}})$.
@@ 1.2   Concavity of constrained optimization (5)
$\geq\sum_{i\neq j}J_{ji}(\lambda)\ln\frac{J_{ji}(\lambda)}{J_{ij}(\lambda)}=\dot{\Sigma}({\boldsymbol{J}}(\lambda)).$ (S10)
Here we used the log-sum inequality [50, Thm. 2.7.1],
$\lambda a\ln\frac{a}{a^{\prime}}+(1-\lambda)b\ln\frac{b}{b^{\prime}}\geq[\lambda a+(1-\lambda)b]\ln\frac{\lambda a+(1-\lambda)b}{\lambda a^{\prime}+(1-\lambda)b^{\prime}}\qquad\qquad\text{for all }a,a^{\prime},b,b^{\prime}\geq 0\,.$ (S11)
1.3   Properties of the unconstrained optimization (6)
1.3.1   Concavity and existence of maximizer
@@
$\mathscr{L}(\bm{p})=\dot{\mathcal{F}}^{b}(\bm{p})+\dot{\mathcal{G}}^{b}(\bm{p})=-\dot{S}^{b}(\bm{p})+\sum_{i}p_{i}\phi_{i},$ (S12)
where $\dot{S}^{b}(\bm{p})=-\sum_{i,j}p_{i}R^{b}_{ji}\ln p_{j}$ is the rate of increase of the Shannon entropy of distribution $\bm{p}$ under the rate matrix $R^{b}$. Here we show that $\mathscr{L}(\bm{p})$ is concave, thus (6) involves the maximization of a concave function. This is equivalent to the minimization of a convex function, making it a so-called “convex optimization problem” which can be solved using standard numerical techniques. We also show the existence of a maximizer
@@
Existence of $\bm{p}^{*}$: The feasible set of the optimization is compact ($n$-dimensional probability simplex) and the objective is continuous, so the maximizer exists by the extreme value theorem. Moreover, the maximum value is finite (see (S87) below).
@@
$\mathscr{L}(\bm{p}(\lambda))\geq\lambda\mathscr{L}(\bm{p})+(1-\lambda)\mathscr{L}(\bm{q}).$ (S13)
Observe that $\sum_{i}p_{i}\phi_{i}$ is linear in $\bm{p}$, thus we can use (S12) to rearrange (S13) as
$\dot{S}^{b}(\bm{p}(\lambda))\leq\lambda\dot{S}^{b}(\bm{p})+(1-\lambda)\dot{S}^{b}(\bm{q}).$ (S14)
To prove (S14), we first rewrite the rate of decrease of the Shannon entropy as
$\dot{S}^{b}(\bm{p})=-\sum_{i,j}p_{i}R^{b}_{ji}\ln p_{j}=-\sum_{i,j}p_{i}R^{b}_{ji}\ln\frac{p_{j}}{p_{i}}=-\sum_{i\neq j}p_{i}R^{b}_{ji}\ln\frac{p_{j}}{p_{i}}=\sum_{i\neq j}p_{i}R^{b}_{ji}\ln\frac{p_{i}}{p_{j}}.$ (S15)
Here we first used that $\sum_{i,j}p_{i}R^{b}_{ji}a_{i}=\sum_{i}a_{i}p_{i}\sum_{j}R^{b}_{ji}=0$ for any $\bm{a}$, then used that the diagonal terms of the sum ($i=j$) vanish. Next, we use (S15) to prove (S14) as
@@
$\leq\sum_{i\neq j}\left[\lambda p_{i}R^{b}_{ji}\ln\frac{p_{i}R^{b}_{ji}}{p_{j}R^{b}_{ji}}+(1-\lambda)q_{i}R^{b}_{ji}\ln\frac{q_{i}R^{b}_{ji}}{q_{j}R^{b}_{ji}}\right]$ (S16)
$=\lambda\sum_{i\neq j}p_{i}R^{b}_{ji}\ln\frac{p_{j}}{p_{i}}+(1-\lambda)\sum_{i\neq j}q_{i}R^{b}_{ji}\ln\frac{q_{j}}{q_{i}}\qquad=\lambda\dot{S}^{b}(\bm{p})+(1-\lambda)\dot{S}^{b}(\bm{q})\,.$
The second line follows from the log-sum inequality (S11).
1.3.2   Strict concavity and uniqueness of maximizer under irreducibility assumption
@@
Strict concavity: The log-sum inequality (S11) is strict when $a/a^{\prime}\neq b/b^{\prime}$ [50, Thm. 2.7.1]. In our case, (S16) is strict for any pair of states $j\neq i$ that have $R^{b}_{ji}>0$ and
$\frac{p_{i}R^{b}_{ji}}{p_{j}R^{b}_{ji}}\neq\frac{q_{i}R^{b}_{ji}}{q_{j}R^{b}_{ji}}.$ (S17)
Such a pair of states must exist, as we now prove by contradiction. Suppose to the contrary that (S17) is an equality for all $i\neq j$ where $R^{b}_{ji}>0$. Consider a walk on the graph defined by the allowed transitions in $R^{b}$ — that is, a sequence of states $i_{0},i_{1},i_{2},\dots$ such that $R^{b}_{i_{k+1}i_{k}}>0$. We would then have that $p_{i_{0}}/p_{i_{1}}=q_{i_{0}}/q_{i_{1}},p_{i_{1}}/p_{i_{2}}=q_{i_{1}}/q_{i_{2}},\dots$, hence also (by multiplication) that $p_{i_{0}}/p_{i_{2}}=q_{i_{0}}/q_{i_{2}},\dots$. Now, since $R^{b}$ is irreducible, any state can be reached from any other, implying that $p_{i}/p_{j}=q_{i}/q_{j}$ for all $i\neq j$, hence $\bm{p}\propto\bm{q}$. Since probabilities are nonnegative and normalized to sum to 1, this can only hold when $\bm{p}=\bm{q}$, contradicting our assumption that $\bm{p}\neq\bm{q}$ in (S16).
To summarize, there must be some $i\neq j$ such that the corresponding term in (S16) is a strict inequality, therefore
$\mathscr{L}(\bm{p}(\lambda))>\lambda\mathscr{L}(\bm{p})+(1-\lambda)\mathscr{L}(\bm{q}).$ (S18)
Uniqueness of maximizer: Suppose that there are two different maximizers, $\mathscr{L}(\bm{p}^{*}_{(1)})=\mathscr{L}(\bm{p}^{*}_{(2)})=\mathscr{G}$. Then, (S18) implies that the convex mixture $\lambda\bm{p}^{*}_{(1)}+(1-\lambda)\bm{p}^{*}_{(2)}$ would achieve a larger value than $\mathscr{G}$, leading to a contradiction.
Full support of $\bm{p}^{*}$: Suppose that the maximizer $\bm{p}^{*}$ did not have full support, so that some states have 0 probability. Then, there must be some pair $i\neq j$ such that $p_{i}^{*}>0$ and $p_{j}^{*}=0$ and $R^{b}_{ji}>0$ (otherwise the graph of allowed transitions would not be strongly connected). But for this pair, $p_{i}^{*}R^{b}_{ji}\ln({p_{j}^{*}}/{p_{i}^{*}})=-\infty$. Plugging into (S15) implies that $\mathscr{L}(\bm{p}^{*})=-\infty$ and contradicting the fact that $\bm{p}^{*}$ is a maximizer. Hence, $\bm{p}^{*}$ must have full support.
1.4   Derivation and achievability of unconstrained optimization (6)

### 12 Mar 2024 v4

@@ Acknowledgements.
We thank members of the Complex Systems Lab, B. Corominas-Murtra, L. Seoane, D. Wolpert, and D. Sowinski for useful discussions. A.K. also thanks Sosuke Ito for support and encouragement. This project has received funding from the European Union’s Horizon 2020 research and innovation programme under the Marie Skłodowska-Curie Grant Agreement No. 101068029. J.P. was supported by the María de Maeztú fellowship MDM-2014-0370-17-2 and Grant No. 62417 from the John Templeton Foundation. The opinions expressed in this publication are those of the authors and do not necessarily reflect the views of the John Templeton Foundation.
@@
Optimization of nonequilibrium free energy harvesting illustrated on bacteriorhodopsin
@@
I Derivations of main results
I.1 Derivation of (8) from LDB and steady-state assumption
@@
$\dot{\mathcal{G}}^{\mathrm{tot}}=\dot{\mathcal{F}}^{b}(\bm{\pi})+\dot{\mathcal{G}}^{b}(\bm{\pi})-\dot{\Sigma}({\boldsymbol{J}}^{c})/\beta.$ (S1.1)
@@
$={\beta}^{-1}\sum_{i,j}p_{i}R^{c}_{ji}[\ln p_{j}+\beta(f_{j}+{g}^{c}_{ji})].$ (S1.2)
The first line follows from definitions made in the main text. Since $R^{c}$ is a rate matrix, $\sum_{i,j}p_{i}R^{c}_{ji}a_{i}=\sum_{i}p_{i}a_{i}\sum_{j}R^{c}_{ji}=0$ for any $\bm{a}$. This allows us to further rewrite (S1.2) as
@@
$\ln(R^{c}_{ji}/R^{c}_{ij})=\beta(f_{i}-f_{j}-{g}^{c}_{ji})\qquad\text{for}\quad R^{c}_{ji}>0.$ (S1.4)
Using the definition of one-way fluxes due to control given a generic distribution $\bm{p}$, ${J^{c}_{ji}}=p_{i}R^{c}_{ji}$, and the Schnackenberg expression of EPR $\dot{\Sigma}$, we rewrite (S1.3) as
˙$\dot{\mathcal{F}}^{c}(\bm{p})+\dot{\mathcal{G}}^{c}(\bm{p})=-{\beta}^{-1}\sum_{i,j}p_{i}R^{c}_{ji}\ln\frac{p_{j}R^{c}_{ji}}{p_{i}R^{c}_{ij}}=-{\beta}^{-1}\sum_{i\neq j}{J^{c}_{ji}}\ln\frac{{J^{c}_{ji}}}{{J^{c}_{ij}}}=-{\beta}^{-1}\dot{\Sigma}({\boldsymbol{J}}^{c})\,.$ (S1.5)
In the main text, this equality is applied at the steady-state distribution $\bm{\pi}$. However, Eq. (S1.5) is valid for all distributions $\bm{p}$ as long as (S1.4) holds.
To complete our derivation of (S1.1), recall that $R\bm{\pi}=(R^{b}+R^{c})\bm{\pi}=0$ by definition of the steady-state $\bm{\pi}$. At the same time, from the definition of ˙$\dot{\mathcal{F}}^{b}(\bm{p})$ in (2), we have
˙$\dot{\mathcal{F}}^{b}(\bm{\pi})+\dot{\mathcal{F}}^{c}(\bm{\pi})=\sum_{j}(f_{j}+{\beta}^{-1}\ln\pi_{j})\sum_{i}(R^{b}_{ji}+R^{c}_{ji})\pi_{i}=0\qquad\implies\qquad\dot{\mathcal{F}}^{b}(\bm{\pi})=-\dot{\mathcal{F}}^{c}(\bm{\pi}).$ (S1.6)
Combining the definition of $\dot{\mathcal{G}}^{\mathrm{tot}}$ with (S1.5) and (S1.6) gives Eq. (S1.1),
$\dot{\mathcal{G}}^{\mathrm{tot}}=\dot{\mathcal{G}}^{b}(\bm{\pi})+\dot{\mathcal{G}}^{c}(\bm{\pi})=\dot{\mathcal{G}}^{b}(\bm{\pi})-\left[\dot{\mathcal{F}}^{c}(\bm{\pi})+{\beta}^{-1}\dot{\Sigma}({\boldsymbol{J}}^{c})\right]=\dot{\mathcal{F}}^{b}(\bm{\pi})+\dot{\mathcal{G}}^{b}(\bm{\pi})-{\beta}^{-1}\dot{\Sigma}({\boldsymbol{J}}^{c})\,.$ (S1.7)
@@
$\dot{\Sigma}({\boldsymbol{J}}^{c})=\frac{1}{2}\sum_{i\neq j}(p_{j}R^{c}_{ij}-p_{i}R^{c}_{ji})\ln\frac{p_{j}R^{c}_{ij}}{p_{i}R^{c}_{ji}}\geq 0,$ (S1.8)
@@
Second, recall our assumption that the control processes do not interact directly with the external source of free energy, but only with the internal reservoir and heat bath. This assumption is formalized in the particular form of the LDB condition (S1.4). It means that for the control processes, the combined “system+internal reservoir” may be treated as a single system coupled only to a heat bath at inverse temperature $\beta$. Hence, the rate of entropy production due to control is simply $\beta$ times the decrease of the combined free energy of the system and internal reservoir, as in (S1.5). Moreover, owing to Second Law, their combined free energy cannot increase under control dynamics [26, 29],
˙$\dot{\mathcal{F}}^{c}(\bm{p})+\dot{\mathcal{G}}^{c}(\bm{p})\leq 0\qquad\text{for all}\quad\bm{p}.$ (S1.9)
I.2 Concavity of constrained optimization (10)
Consider the objective ˙$\dot{\mathcal{F}}^{b}(\bm{p})+\dot{\mathcal{G}}^{b}(\bm{p})-\dot{\Sigma}({\boldsymbol{J}})/\beta$ in (10) in the main text. In Sec. I.3 below, we show that the term ˙$\dot{\mathcal{F}}^{b}(\bm{p})+\dot{\mathcal{G}}^{b}(\bm{p})$ is concave in $\bm{p}$. Here we show that $\dot{\Sigma}({\boldsymbol{J}})$ is convex in ${\boldsymbol{J}}$. This shows that the overall objective is concave in the pair $(\bm{p},{\boldsymbol{J}})$.
@@ I.2 Concavity of constrained optimization (10)
$\geq\sum_{i\neq j}J_{ji}(\lambda)\ln\frac{J_{ji}(\lambda)}{J_{ij}(\lambda)}=\dot{\Sigma}({\boldsymbol{J}}(\lambda)).$ (S1.10)
Here we used the log-sum inequality [74, Thm. 2.7.1],
$\lambda a\ln\frac{a}{a^{\prime}}+(1-\lambda)b\ln\frac{b}{b^{\prime}}\geq[\lambda a+(1-\lambda)b]\ln\frac{\lambda a+(1-\lambda)b}{\lambda a^{\prime}+(1-\lambda)b^{\prime}}\qquad\qquad\text{for all }a,a^{\prime},b,b^{\prime}\geq 0\,.$ (S1.11)
I.3 Properties of the unconstrained optimization (11)
I.3.1 Concavity and existence of maximizer
@@
$\mathscr{L}(\bm{p})=\dot{\mathcal{F}}^{b}(\bm{p})+\dot{\mathcal{G}}^{b}(\bm{p})=-\dot{S}^{b}(\bm{p})+\sum_{i}p_{i}\phi_{i},$ (S1.12)
where $\dot{S}^{b}(\bm{p})=-\sum_{i,j}p_{i}R^{b}_{ji}\ln p_{j}$ is the rate of increase of the Shannon entropy of distribution $\bm{p}$ under the rate matrix $R^{b}$. Here we show that $\mathscr{L}(\bm{p})$ is concave, thus (11) involves the maximization of a concave function. This is equivalent to the minimization of a convex function, making it convex optimization problem which can be solved using standard numerical techniques. We also show the existence of a maximizer
@@
Existence of $\bm{p}^{*}$: The feasible set is compact ($n$-dimensional probability simplex) and the objective is continuous, so the maximizer exists by the extreme value theorem. Also, the maximum value is finite (see (S3.86)).
@@
$\mathscr{L}(\bm{p}(\lambda))\geq\lambda\mathscr{L}(\bm{p})+(1-\lambda)\mathscr{L}(\bm{q}).$ (S1.13)
Observe that $\sum_{i}p_{i}\phi_{i}$ is linear in $\bm{p}$, thus we can use (S1.12) to rearrange (S1.13) as
$\dot{S}^{b}(\bm{p}(\lambda))\leq\lambda\dot{S}^{b}(\bm{p})+(1-\lambda)\dot{S}^{b}(\bm{q}).$ (S1.14)
To prove (S1.14), we first rewrite the rate of decrease of the Shannon entropy as
$\dot{S}^{b}(\bm{p})=-\sum_{i,j}p_{i}R^{b}_{ji}\ln p_{j}=-\sum_{i,j}p_{i}R^{b}_{ji}\ln\frac{p_{j}}{p_{i}}=-\sum_{i\neq j}p_{i}R^{b}_{ji}\ln\frac{p_{j}}{p_{i}}=\sum_{i\neq j}p_{i}R^{b}_{ji}\ln\frac{p_{i}}{p_{j}}.$ (S1.15)
Here we first used that $\sum_{i,j}p_{i}R^{b}_{ji}a_{i}=\sum_{i}a_{i}p_{i}\sum_{j}R^{b}_{ji}=0$ for any $\bm{a}$, then used that the diagonal terms of the sum ($i=j$) vanish. Next, we use (S1.15) to prove (S1.14) as
@@
$\leq\sum_{i\neq j}\left[\lambda p_{i}R^{b}_{ji}\ln\frac{p_{i}R^{b}_{ji}}{p_{j}R^{b}_{ji}}+(1-\lambda)q_{i}R^{b}_{ji}\ln\frac{q_{i}R^{b}_{ji}}{q_{j}R^{b}_{ji}}\right]$ (S1.16)
$=\lambda\sum_{i\neq j}p_{i}R^{b}_{ji}\ln\frac{p_{j}}{p_{i}}+(1-\lambda)\sum_{i\neq j}q_{i}R^{b}_{ji}\ln\frac{q_{j}}{q_{i}}$
$=\lambda\dot{S}^{b}(\bm{p})+(1-\lambda)\dot{S}^{b}(\bm{q})\,.$
The second line follows from the log-sum inequality (S1.11).
I.3.2 Strict concavity and uniqueness of maximizer under irreducibility assumption
@@
Strict concavity: The log-sum inequality (S1.11) is strict when $a/a^{\prime}\neq b/b^{\prime}$ [74, Thm. 2.7.1]. In our case, (S1.16) is strict for any pair of states $j\neq i$ that have $R^{b}_{ji}>0$ and
$\frac{p_{i}R^{b}_{ji}}{p_{j}R^{b}_{ji}}\neq\frac{q_{i}R^{b}_{ji}}{q_{j}R^{b}_{ji}}.$ (S1.17)
Such a pair of states must exist, as we now prove by contradiction. Suppose to the contrary that (S1.17) is an equality for all $i\neq j$ where $R^{b}_{ji}>0$. Consider a walk on the graph defined by the allowed transitions in $R^{b}$ — that is, a sequence of states $i_{0},i_{1},i_{2},\dots$ such that $R^{b}_{i_{k+1}i_{k}}>0$. We would then have that $p_{i_{0}}/p_{i_{1}}=q_{i_{0}}/q_{i_{1}},p_{i_{1}}/p_{i_{2}}=q_{i_{1}}/q_{i_{2}},\dots$, hence also (by multiplication) that $p_{i_{0}}/p_{i_{2}}=q_{i_{0}}/q_{i_{2}},\dots$. Now, since $R^{b}$ is irreducible, any state can be reached from any other, implying that $p_{i}/p_{j}=q_{i}/q_{j}$ for all $i\neq j$, hence $\bm{p}\propto\bm{q}$. Since probabilities are nonnegative and normalized to sum to 1, this can only hold when $\bm{p}=\bm{q}$, contradicting our assumption that $\bm{p}\neq\bm{q}$ in (S1.16).
To summarize, there must be some $i\neq j$ such that the corresponding term in (S1.16) is a strict inequality, therefore
$\mathscr{L}(\bm{p}(\lambda))>\lambda\mathscr{L}(\bm{p})+(1-\lambda)\mathscr{L}(\bm{q}).$ (S1.18)
Uniqueness of maximizer: Suppose that there are two different maximizers, $\mathscr{L}(\bm{p}^{*}_{(1)})=\mathscr{L}(\bm{p}^{*}_{(2)})=\mathscr{G}$. Then, (S1.18) implies that the convex mixture $\lambda\bm{p}^{*}_{(1)}+(1-\lambda)\bm{p}^{*}_{(2)}$ would achieve a larger value than $\mathscr{G}$, leading to a contradiction.
Full support of $\bm{p}^{*}$: Suppose that the maximizer $\bm{p}^{*}$ did not have full support, so that some states have 0 probability. Then, there must be some pair $i\neq j$ such that $p_{i}^{*}>0$ and $p_{j}^{*}=0$ and $R^{b}_{ji}>0$ (otherwise the graph of allowed transitions would not be strongly connected). But for this pair, $p_{i}^{*}R^{b}_{ji}\ln({p_{j}^{*}}/{p_{i}^{*}})=-\infty$. Plugging into (S1.15) implies that $\mathscr{L}(\bm{p}^{*})=-\infty$ and contradicting the fact that $\bm{p}^{*}$ is a maximizer. Hence, $\bm{p}^{*}$ must have full support.
I.4 Derivation and achievability of unconstrained optimization (11)